# Fine-Grained Alignment and Noise Refinement for Compositional Text-to-Image Generation

### v1

@@ 1 Introduction
Recent advancements in diffusion-based text-to-image (T2I) models [34, 29, 36, 33, 4] have significantly improved the generation of high-quality and diverse images from textual prompts. However, these models often fail to precisely capture the intended meaning, resulting in inconsistencies between the generated images and the original prompt. Recent studies [18, 17, 26, 12] highlight key failure modes in text-to-image generation, including entity missing, attribute binding errors, and incorrect relationships. In response, training-free approaches have been introduced to tackle the challenge of compositionality. The first approach uses spatial layouts or bounding boxes to guide generation [22, 50, 39, 52]. While these methods enhance spatial coherence by enforcing structured layouts, they often struggle with visual realism and maintaining aesthetic quality [50, 51]. The second branch utilizes Large Language Models (LLMs) to decompose complex tasks into manageable subtasks [24, 42, 21, 45]. The effectiveness of these methods heavily depends on the capabilities of the underlying LLMs and the quality of prompt engineering [42, 51]. The third branch focuses on optimizing attention maps [7, 26, 23, 37, 2, 12]. Although these methods typically address one or two specific failure modes, they fail to comprehensively resolve all of them. For instance, A-STAR [2], A&E [7], and InitNO [12] primarily tackle entity mixing and missing issues, while Divide-and-Bind [23] and CONFORM [26] go further by addressing attribute-binding errors as well. Yet, a unified approach that holistically addresses compositionality challenges remains an open problem. We follow the attention-map approach and propose a novel training-free method that defines objectives to arrange the scene based on the entities and their relations mentioned in the prompt. To achieve this, we define four loss functions: (1) object missing loss, which encourages the presence of each object in the attention map; (2) object mixing loss, which ensures that each object occupies a distinct space in the attention map; (3) attribute-binding loss, which guides the attention maps of an entity and its attributes to cover the same area, thereby forcing an entity to have its own attributes; and (4) spatial relation loss, which constrains the attention maps of entities to align with the spatial relationships described in the corresponding text components.
@@
Moreover, we take an additional step by designing a feedback-driven system that refines the initial noise of the first generation stage. This system incorporates a verifier that evaluates the generated image and provides feedback on inconsistencies. Our approach leverages feedback signal to identify and correct faulty regions within the initial noise. Specifically, we expand the initial noise into attention maps, refine the faulty ones based on the feedback, by adjusting the noise, thus providing a better starting point for the generation process. Then, the second phase proceeds by applying our loss functions during the generation process. To the best of our knowledge, this form of selective refinement for initial noise has not been explored in previous studies. In contrast to our approach, existing methods such as InitNO [12], CoCoNO [38], and ReNO [9] regenerate the entire noise without distinguishing which parts are faulty. For example, in Fig. 1, given the prompt ”a dog is sitting next to its owner”, the first generated image suffers from an entity missing issue, where the ”owner” is not correctly represented. The verifier identifies this inconsistency and provides feedback. Consequently, our method automatically activates the entity missing loss and applies it to the owner’s attention map in the initial noise, and then proceeds with the generation process guided by our losses. Also, this form of initial noise refinement can be generalized to handle multiple objects and failure modes simultaneously. We use DA-Score [37] as our primary verifier, enhancing its fine-grained question extraction with additional coarse-grained questions. These questions are grouped by failure mode, allowing for precise error detection. The primary contributions of our work are outlined as follows:
@@ 3.1 Fine-grained Alignment by EAR Loss
$\begin{split}\mathcal{L}_{\text{spatial}}=&\!\!\!\!\!\!\!\!\!\!\!\!\sum_{(e_{1},r,e_{2})\in G_{\text{relations}}}\\ \Big{[}&\mathbb{I}_{x}(r)\cdot\sigma\Big{(}dir(r)\cdot\big{(}E_{x}(A_{e_{2}}^{t})-E_{x}(A_{e_{1}}^{t})\big{)}\Big{)}\\ +&\mathbb{I}_{y}(r)\cdot\sigma\Big{(}dir(r)\cdot\big{(}E_{y}(A_{e_{2}}^{t})-E_{y}(A_{e_{1}}^{t})\big{)}\Big{)}\Big{]},\end{split}$ (5)
where $\mathbb{I(.)}$ and $dir(.)$ represent the indicator and direction functions, respectively. Both are formally defined in 2. To illustrate the above formula, in Fig. 2, we consider the triple (cat, left, frog). The spatial loss formula for this case is expressed as:
@@
Latent Update: During the first 25 steps of the diffusion process, the latent space at timestep $t$, denoted as $Z_{t}$, is denoised using the U-Net [35], yielding $Z^{\prime}_{t-1}$. It is then refined using the following update rule:
@@ 3.2 Fine-grained Initial Noise Refinement
Misalignment Detection: In this stage, we assess the quality of the first generated image and identify potential flaws. To achieve this, we select a verifier, such as DA-Score [37], and adapt it to provide fine-grained feedback on specific types of issues present in the image. An example of this misalignment detection process is illustrated on the left side of Fig. 3, Further details on misalignment detection and verifier adaptation are provided in Supp. 8.
@@
$\begin{split}\mathcal{L}_{correction}&=\\ \mathbb{I}_{\text{entity}}&\mathcal{L}^{f}_{\text{entity}}+\mathbb{I}_{\text{attribute}}\mathcal{L}^{f}_{\text{attribute}}+\mathbb{I}_{\text{spatial}}\mathcal{L}^{f}_{\text{spatial}},\end{split}$ (9)
@@
10:     $L_{\text{preservation}}\leftarrow\text{Compute }L_{\text{preservation}}\text{ from }$
11:         proper_entities, $A^{init}\text{, and }A^{\text{ref}}$vartriangleright Eq. 10
12:     $L_{\text{refinement}}\leftarrow L_{\text{preservation}}+L_{\text{correction}}$vartriangleright Eq. 11
13:     $Z_{T}\leftarrow Z_{T}-\alpha\nabla_{Z_{T}}L_{\text{refinement}}$vartriangleright Eq. 12
@@
15:  end while
16:  $\mathcal{I}\leftarrow G(Z_{T},\mathcal{P})$
17:  return  $\mathcal{I}$
4 Experiments
@@ 4.1 Experimental Settings
Datasets. The evaluation is performed on T2I-CompBench [18] and HRS [3]. T2I-CompBench assesses complex compositional generation, while HRS is used to evaluate our model’s capability in multi-entity generation, we use the HRS dataset’s color 3-entity composition category. More information on the datasets is available in the Supp. 9.1. Baselines. We compare ourselves with four categories of methods: (1) base models including stable diffusion v1.4, v2 [34] and XL [30]; (2) approaches focused on latent space optimization such as A&E [7], CONFORM [26], D&B [23] and A-STAR [2]; (3) Evaluate&Refine [37], which utilizes Visual Question Answering (VQA) feedback; and (4) InitNO [12], which iteratively alters the initial noise. In experiments, all baselines, excluding the Stable Diffusion family, are based on Stable Diffusion v1.4.
@@ 4.2 Quantitative Results
We quantitatively evaluate our approach using multiple verifiers. Table 1 and 2 present the performance of all methods based on the DA-Score [37] and TIFA-Score [17], respectively. Although some of these models generate images iteratively, our one-stage generation method outperformed the others in most categories. Table 3 presents our results on the HRS dataset, focusing on three-entity prompts from the color category. This experiment highlights the effectiveness of our method in handling multiple entities. Our fine-grained noise refinement significantly enhances the results, achieving improvements of up to 7% on the DA-Score and 4% on the Tifa-Score, respectively. This underscores the effectiveness of our fine-grained noise refinement method in handling multi-object prompts. To ensure a reliable evaluation of generative models on spatial relationships, we employ the VISOR Score [11], which precisely evaluates spatial relationships. To perform this experiment, we require a larger set of prompts. Therefore, we arbitrarily select 150 spatial prompts from the T2I-CompBench [18] training dataset. As shown in Figure 4(a), integrating our spatial loss with other losses resulted in a substantial 25% boost in the VISOR Score, outperforming other baselines. Additional quantitative results are provided in the Supp. 9.4
@@
Table 1: Performance Comparison. Evaluation of various models on the DA-Score metric across different categories using the T2I-CompBench benchmark. The term ”iter” in this table indicates the number of times the image is generated using that specific method.
@@
Table 3: Performance Comparison. Evaluation of various models using the HRS benchmark on the DA-Score and Tifa-Score metrics. The term ”iter” in this table indicates the number of times the image is generated using that specific method.
@@
User Study. We conducted a user study with 15 volunteers to evaluate image-text alignment, comparing our method against three baseline models: CONFORM [26], InitNO [12], and Evaluate&Refine [37]. Using 25 prompts from T2I-CompBench, volunteers selected the best-matching image per prompt. Our method outperformed the baselines by 24%, as shown in Figure 4(b). More detail is reported in Supp. 9.2.2
@@
Inference-time and Diversity Comparison Table 4 presents a comparison of inference time and FID [15] scores across various baselines. We randomly selected 50 sample prompts from T2I-Combench [18] to evaluate and benchmark the performance of our method against existing approaches. Our one stage method achieves an average inference time of $1$ 0.15 seconds, closely matching SDv1 (11.10 ± 0.13 seconds) while significantly outperforming computationally heavier methods such as DA-Score (23.26s) and InitNO (32.56s). Moreover, fine-grained initial noise refinement increases the inference time to 22.34 ± 0.50 seconds but improves the FID score to 190.47, demonstrating a trade-off between efficiency and generation quality. Compared to other baselines, our method strikes a balance between fast inference and competitive FID performance. Additional comparisons on diversity and aesthetics can be found in Supp. 9.3.2.
@@
4.3 Qualitative Results.
Figure 5 presents a qualitative comparison of our proposed method against existing models using similar prompts, while Figure 6 compares our approach with the state-of-the-art Stable Diffusion model, SD-XL [30]. In Figure 5, our method demonstrates superior attribute binding, particularly evident in the shapes depicted in ”An oval sink and a rectangular mirror.”. Also, regarding spatial relation, all baseline models fail to correctly interpret the concept of right, as seen in ”A boy on the right of a chicken.”. In Figure 6, our method, based on SD v1.4, outperforms SD-XL [30] across all categories, including entity missing, attribute binding, and spatial relation. Additional qualitative results are available in Supp. 9.5
Attention maps function as object detection. In Fig 7, we demonstrate that our last step cross attention maps can be sufficiently rich to serve as object detectors. Comparing our method with Evaluate&Refine [37], another soft multi-iteration approach, we show that our method is both faithful and location-aware, effectively arranging scenes in a coherent manner when prompts contain multiple entities. Further details and visualization can be found in Supp. 9.6.

### v2

@@ 1 Introduction
Recent advancements in diffusion-based text-to-image (T2I) models (Rombach et al., 2022; Nichol et al., 2021; Saharia et al., 2022; Ramesh et al., 2022; Balaji et al., 2022) have significantly improved the generation of high-quality and diverse images from textual prompts. However, these models often fail to precisely capture the intended meaning, resulting in inconsistencies between the generated images and the original prompt. Recent studies (Huang et al., 2023a; Hu et al., 2023; Meral et al., 2024; Guo et al., 2024) highlight key failure modes in text-to-image generation, including entity missing, attribute binding errors, and incorrect relationships. In response, training-free approaches have been introduced to tackle the challenge of compositionality. The first approach uses spatial layouts or bounding boxes to guide generation (Li et al., 2023; Zhang et al., 2024b; Wang et al., 2024; Zheng et al., 2023). While these methods enhance spatial coherence by enforcing structured layouts, they often struggle with visual realism and maintaining aesthetic quality (Zhang et al., 2024b; 2025). The second branch utilizes Large Language Models (LLMs) to decompose complex tasks into manageable subtasks (Lian et al., 2023; Yang et al., 2024; Li et al., 2024a; Ye et al., 2024). The effectiveness of these methods heavily depends on the capabilities of the underlying LLMs and the quality of prompt engineering (Yang et al., 2024; Zhang et al., 2025). The third branch focuses on optimizing attention maps (Chefer et al., 2023b; Meral et al., 2024; Li et al., 2024b; Singh & Zheng, 2023; Agarwal et al., 2023; Guo et al., 2024). Although these methods typically address one or two specific failure modes, they fail to comprehensively resolve all of them. For instance, A-STAR (Agarwal et al., 2023), A&E (Chefer et al., 2023b), and InitNO (Guo et al., 2024) primarily tackle entity mixing and missing issues, while Divide-and-Bind (Li et al., 2024b) and CONFORM (Meral et al., 2024) go further by addressing attribute-binding errors as well. Yet, a unified approach that holistically addresses compositionality challenges remains an open problem. We follow the attention-map approach and propose a novel training-free method that defines objectives to arrange the scene based on the entities and their relations mentioned in the prompt. To achieve this, we define four loss functions: (1) object missing loss, which encourages the presence of each object in the attention map; (2) object mixing loss, which ensures that each object occupies a distinct space in the attention map; (3) attribute-binding loss, which guides the attention maps of an entity and its attributes to cover the same area, thereby forcing an entity to have its own attributes; and (4) spatial relation loss, which constrains the attention maps of entities to align with the spatial relationships described in the corresponding text components.
@@
Moreover, we take an additional step by designing a feedback-driven system that refines the initial noise of the first generation stage. This system incorporates a verifier that evaluates the generated image and provides feedback on inconsistencies. Our approach leverages feedback signal to identify and correct faulty regions within the initial noise. Specifically, we expand the initial noise into attention maps, refine the faulty ones based on the feedback, by adjusting the noise, thus providing a better starting point for the generation process. Then, the second phase proceeds by applying our loss functions during the generation process. To the best of our knowledge, this form of selective refinement for initial noise has not been explored in previous studies. In contrast to our approach, existing methods such as InitNO (Guo et al., 2024), CoCoNO (Sundaram et al., 2024), and ReNO (Eyring et al., 2024) regenerate the entire noise without distinguishing which parts are faulty. For example, in Fig. 1, given the prompt "a dog is sitting next to its owner", the first generated image suffers from an entity missing issue, where the "owner" is not correctly represented. The verifier identifies this inconsistency and provides feedback. Consequently, our method automatically activates the entity missing loss and applies it to the owner’s attention map in the initial noise, and then proceeds with the generation process guided by our losses. Also, this form of initial noise refinement can be generalized to handle multiple objects and failure modes simultaneously. For instance, in Fig. 1, given the prompt "The leather belt and fluffy hat hang on the metallic hook by the glass door", the initial generated image not only omits the hat but also fails to render the door with a glass-like appearance. Our approach leverages these feedback signals to correct both issues. while preserving correctly generated elements in the noise, such as the belt. We use DA-Score (Singh & Zheng, 2023) as our primary verifier, enhancing its fine-grained question extraction with additional coarse-grained questions. These questions are grouped by failure mode, allowing for precise error detection. The primary contributions of our work are outlined as follows:
@@
2 Related Works
Before the emergence of diffusion-based models, various research directions aimed to achieve realistic and high-quality image generation, both conditional and unconditional, through approaches like generative adversarial networks (GANs) (Kang et al., 2023; Xu et al., 2018; Ye et al., 2021; Zhang et al., 2021; Zhu et al., 2019; Zhang et al., 2017) and autoregressive models (Chang et al., 2023; Ding et al., 2021; Ramesh et al., 2021; Yu et al., 2022). Recently, diffusion-based models have shown remarkable performance in conditional text-to-image generation (Rombach et al., 2022; Nichol et al., 2021; Saharia et al., 2022; Ramesh et al., 2022; Balaji et al., 2022). While having a good performance, they often have alignment problems when generating images. We discuss previous studies that attempt to alleviate alignment problems under three categories.
Training-Based Compositional Improvement. A group of studies have focused on overcoming compositional challenges in text-conditional diffusion-based models to improve alignment between text prompts and generated images. Some approaches address this through training-based methods (Li et al., 2023; Yang et al., 2023; Mou et al., 2024; Zhang et al., 2023; Huang et al., 2023b; Zhang et al., 2025; Eyring et al., 2024). For instance, T2I-Adapter (Mou et al., 2024) and ControlNet (Zhang et al., 2023) target the control of semantic structures by specifying high-level features. ReCo (Yang et al., 2023) refines spatial awareness using adapters on top of the diffusion models, while GLIGEN (Li et al., 2023) integrates grounding information into newly trainable layers. Composer (Huang et al., 2023b) decomposes images into key factors, training a diffusion model with these factors as conditions to recompose the input. Alternatively, Itercomp (Zhang et al., 2025) utilizes iterative feedback learning to enhance the compositional generation, and ReNo (Eyring et al., 2024) adopts reward-based noise optimization for improved alignment. While being effective, adopting these methods comes with the cost of additional training time and resources.
Training-Free Compositional Improvement. To mitigate compositional problems, some methods leverage training-free methods optimizing latent or attention maps during inference (Liu et al., 2022; Feng et al., 2022; Chefer et al., 2023a; Agarwal et al., 2023; Meral et al., 2024; Guo et al., 2024; Marioriyad et al., 2024; Ma et al., 2025; Li et al., 2025; Yu & Gao, 2025). For instance, Composable Diffusion (Liu et al., 2022) computes separate denoising latent for each statement and combines them with a score function, or Structure Diffusion (Feng et al., 2022) manipulates the cross-attention maps guided by hierarchically extracted structures from text. Attend-and-Excite proposed a novel loss based on cross-attention maps to reduce missing objects. Similarly, A-STAR (Agarwal et al., 2023) optimizes latent representations by employing segregation and retention losses to minimize cross-attention overlap between different concepts, leading the objects not to be overlooked while preserving cross-attention information across all concepts. Also, Divide-and-Bind (Li et al., 2024b) addresses the problem of entity missing by leveraging an attendance objective. Additionally, it enhances attribute association through the incorporation of a binding objective. Similarly, CONFORM (Meral et al., 2024) addresses entity missing and attribute binding issues by employing a contrastive objective. This approach brings each entity and its attribute’s attention map closer together while disentangling the attention maps of different entities and their respective attributes. In another approach (Guo et al., 2024), the initial noise is optimized to ensure it lies within a valid space for generating images aligned with the input prompt. Another approach involves utilizing evaluation feedback to refine the image after its generation. The Evaluate-and-Refine method (Singh & Zheng, 2023) employs iterative VQA feedback to adjust the weighting of CLIP’s phrase embeddings. Additionally, it integrates the Attend-and-Excite method to address the issue of entity missing effectively. However, these approaches are restricted in handling compositional problems such as spatial relationships, entity missing, and attribute binding in an integrated manner while utilizing textual constraints simultaneously enables us to arrange the scene more effectively.
Leveraging LLMs for Accurate Alignment in Text-to-Image Models. Recent works have leveraged layout-based approaches to enhance compositional alignment in images generated by text-to-image models, using spatial layouts or bounding boxes (Lian et al., 2023; Zhang et al., 2024b; Li et al., 2023; Wang et al., 2024; Zheng et al., 2023). Although these methods enhance spatial awareness, they face challenges in achieving realistic image generation, particularly in creating non-spatial relationships and maintaining aesthetic quality (Zhang et al., 2024b; 2025). Other studies utilizing LLMs aim to break down the generation of an aligned image by converting the prompt into simpler subtasks, with the planning of these subtasks throughout the generation process (Lian et al., 2023; Ye et al., 2024; Li et al., 2024a; Yang et al., 2024; Park et al., 2025). However, these models often struggle to produce accurate results due to the inherent complexity of LLM outputs (Yang et al., 2024; Zhang et al., 2025). Additionally, since these methods also extract spatial layouts or bounding boxes, they encounter challenges similar to those faced by layout-based methods.
3 Preliminaries
Stable Diffusion. We apply our alignment and refinement method on the Stable Diffusion model (Rombach et al., 2022). In SD, an encoder $\mathcal{E}$ is trained to map an image $x\in\mathcal{X}$ to a latent representation $z=\mathcal{E}(x)$. The latent code $z$ is then given to a decoder model to reconstruct the input image such that $\mathcal{D}(\mathcal{E}(x))\approx x$. After training the autoencoder, a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) is trained over the latent space of the autoencoder. The UNet (Ronneberger et al., 2015) denoising model $\epsilon_{\theta}$ learns to denoise an input latent $z_{t}$ at each timestep $t$ (where $z_{t}$ results from adding $\epsilon$ noise gradually over $t$ timesteps to the original latent $z_{0}$ during training). The denoising objective which intends to learn $\epsilon_{\theta}$ in order to predict the noise $\epsilon$ is given by:
The denoising process of Stable Diffusion is conditioned on the embedding of text information $L(\mathcal{P})$. In Stable Diffusion, this embedding is the output of the CLIP (Radford et al., 2021) text encoder. At inference time, a random latent $z_{t}$ is sampled from $\mathcal{N}(0,I)$, and the trained UNet denoising model $\epsilon_{\theta}$ outputs a denoised latent $z_{0}$. We obtain the reconstructed image by feeding $z_{0}$ to the decoder $\mathcal{D}$.
The cross-attention map has the shape of $h\times w\times l$ where $h$ and $w$ are feature map resolution and $l$ is the number of text tokens. We set feature map resolution to $16\times 16$ as it empirically contains the most semantically meaningful attention maps (Hertz et al., 2022).
@@ 5.1 Fine-grained Alignment by EAR Loss
$\mathcal{L}_{\text{spatial}}=\sum_{(e_{1},r,e_{2})\in G_{\text{relations}}}\Big{[}\mathbb{I}_{x}(r)\cdot\sigma\big{(}dir(r)\cdot\big{(}E_{x}(A_{e_{2}}^{t})-E_{x}(A_{e_{1}}^{t})\big{)}\big{)}+\mathbb{I}_{y}(r)\cdot\sigma\big{(}dir(r)\cdot\big{(}E_{y}(A_{e_{2}}^{t})-E_{y}(A_{e_{1}}^{t})\big{)}\big{)}\Big{]}$ (6)
where $\mathbb{I(.)}$ and $dir(.)$ represent the indicator and direction functions, respectively. Both are formally defined in 4. To illustrate the above formula, in Fig. 2, we consider the triple (cat, left, frog). The spatial loss formula for this case is expressed as:
@@
Latent Update: During the first 25 steps of the diffusion process, the latent space at timestep $t$, denoted as $Z_{t}$, is denoised using the U-Net (Ronneberger et al., 2015), yielding $Z^{\prime}_{t-1}$. It is then refined using the following update rule:
@@ 5.2 Fine-grained Initial Noise Refinement
Misalignment Detection: In this stage, we assess the quality of the first generated image and identify potential flaws. To achieve this, we select a verifier, such as DA-Score (Singh & Zheng, 2023), and adapt it to provide fine-grained feedback on specific types of issues present in the image. An example of this misalignment detection process is illustrated on the left side of Fig. 3, Further details on misalignment detection and verifier adaptation are provided in Supp. B.
@@
$\mathcal{L}_{correction}=\mathbb{I}_{\text{entity}}\mathcal{L}^{f}_{\text{entity}}+\mathbb{I}_{\text{attribute}}\mathcal{L}^{f}_{\text{attribute}}+\mathbb{I}_{\text{spatial}}\mathcal{L}^{f}_{\text{spatial}},$ (10)
@@
10:     $L_{\text{preservation}}\leftarrow\text{Compute }L_{\text{preservation}}\text{ from }$ proper_entities, $A^{init}\text{, and }A^{\text{ref}}$vartriangleright Eq. 11
11:     $L_{\text{refinement}}\leftarrow L_{\text{preservation}}+L_{\text{correction}}$vartriangleright Eq. 12
12:     $Z_{T}\leftarrow Z_{T}-\alpha\nabla_{Z_{T}}L_{\text{refinement}}$vartriangleright Eq. 13
@@
14:  end while
15:  $\mathcal{I}\leftarrow G(Z_{T},\mathcal{P})$
16:  return  $\mathcal{I}$
6 Results
@@ 6.1 Experimental Settings
Datasets. The evaluation is performed on T2I-CompBench (Huang et al., 2023a) and HRS (Bakr et al., 2023). T2I-CompBench assesses complex compositional generation, while HRS is used to evaluate our model’s capability in multi-entity generation, we use the HRS dataset’s color 3-entity composition category. More information on the datasets is available in the Supp. C.1.
Baselines. We compare ourselves with four categories of methods: (1) base models including stable diffusion v1.4, v2 (Rombach et al., 2022) and XL (Podell et al., 2023); (2) approaches focused on latent space optimization such as A&E (Chefer et al., 2023b), CONFORM (Meral et al., 2024), D&B (Li et al., 2024b) and A-STAR (Agarwal et al., 2023); (3) Evaluate&Refine (Singh & Zheng, 2023), which utilizes Visual Question Answering (VQA) feedback; and (4) InitNO (Guo et al., 2024), which iteratively alters the initial noise. In experiments, all baselines, excluding the Stable Diffusion family, are based on Stable Diffusion v1.4.
@@ 6.2 Quantitative Results
We quantitatively evaluate our approach using multiple verifiers. Table 1 and 2 present the performance of all methods based on the DA-Score (Singh & Zheng, 2023) and TIFA-Score (Hu et al., 2023), respectively. Although some of these models generate images iteratively, our one-stage generation method outperformed the others in most categories. Table 3 presents our results on the HRS dataset, focusing on three-entity prompts from the color category. This experiment highlights the effectiveness of our method in handling multiple entities. Our fine-grained noise refinement significantly enhances the results, achieving improvements of up to 7% on the DA-Score and 4% on the Tifa-Score, respectively. This underscores the effectiveness of our fine-grained noise refinement method in handling multi-object prompts. To ensure a reliable evaluation of generative models on spatial relationships, we employ the VISOR Score (Gokhale et al., 2022), which precisely evaluates spatial relationships. To perform this experiment, we require a larger set of prompts. Therefore, we arbitrarily select 150 spatial prompts from the T2I-CompBench (Huang et al., 2023a) training dataset. As shown in Figure 4(a), integrating our spatial loss with other losses resulted in a substantial 25% boost in the VISOR Score, outperforming other baselines. Additional quantitative results are provided in the Supp. C.4
Table 1: Performance Comparison. Evaluation of various models on the DA-Score metric across different categories using the T2I-CompBench benchmark. The term "iter" in this table indicates the number of times the image is generated using that specific method.
@@
Table 3: Performance Comparison. Evaluation of various models using the HRS benchmark on the DA-Score and Tifa-Score metrics. The term "iter" in this table indicates the number of times the image is generated using that specific method.
@@
User Study. We conducted a user study with 15 volunteers to evaluate image-text alignment, comparing our method against three baseline models: CONFORM (Meral et al., 2024), InitNO (Guo et al., 2024), and Evaluate&Refine (Singh & Zheng, 2023). Using 25 prompts from T2I-CompBench, volunteers selected the best-matching image per prompt. Our method outperformed the baselines by 24%, as shown in Figure 4(b). More detail is reported in Supp. C.2.2
@@
Inference-time and Diversity Comparison Table 5 presents a comparison of inference time and FID (Heusel et al., 2017) scores across various baselines. We randomly selected 50 sample prompts from T2I-Combench (Huang et al., 2023a) to evaluate and benchmark the performance of our method against existing approaches. Our one stage method achieves an average inference time of $1$ 0.15 seconds, closely matching SDv1 (11.10 ± 0.13 seconds) while significantly outperforming computationally heavier methods such as DA-Score (23.26s) and InitNO (32.56s). Moreover, fine-grained initial noise refinement increases the inference time to 22.34 ± 0.50 seconds but improves the FID score to 190.47, demonstrating a trade-off between efficiency and generation quality. Compared to other baselines, our method strikes a balance between fast inference and competitive FID performance. Additional comparisons on diversity and aesthetics can be found in Supp. C.3.2.
Distinction from Test-Time Sampling Methods We emphasize that our method is different from Test-Time Sampling (TTS) approaches (Ma et al., 2025). While TTS methods typically perform multiple forward passes to select an optimal initial noise for generation, our approach follows a refinement-based strategy. Specifically, we start from a single randomly sampled noise and apply up to two refinement stages to improve it as a better starting point for compositional generation. This makes our task inherently more challenging than standard TTS methods, as not all initial noise samples can be effectively refined. To ensure a fair and comprehensive evaluation, we report results for $N=1,2,3$ refinement steps in Table 4. It is worth mentioning that parallel TTS methods like repeated sampling can also be combined with our approach.
Table 4: Evaluation of various models across multiple categories on the T2I-CompBench benchmark.
Model Shape Color Texture Spatial Relation Complex
Best-of-N (N=1) 64 62 66 70 62
Best-of-N (N=2) 71 74 75 76 69
Best-of-N (N=3) 72 78 77 77 70
Ours + FR 72 80 81 81 72
@@
6.3 Qualitative Results
Figure 5 presents a qualitative comparison of our proposed method against existing models using similar prompts, while Figure 6 compares our approach with the state-of-the-art Stable Diffusion model, SD-XL (Podell et al., 2023). In Figure 5, our method demonstrates superior attribute binding, particularly evident in the shapes depicted in "An oval sink and a rectangular mirror.". Also, regarding spatial relation, all baseline models fail to correctly interpret the concept of right, as seen in "A boy on the right of a chicken.". In Figure 6, our method, based on SD v1.4, outperforms SD-XL (Podell et al., 2023) across all categories, including entity missing, attribute binding, and spatial relation. Additional qualitative results are available in Supp. C.5
Attention maps function as object detection. In Fig 7, we demonstrate that our last step cross attention maps can be sufficiently rich to serve as object detectors. Comparing our method with Evaluate&Refine (Singh & Zheng, 2023), another soft multi-iteration approach, we show that our method is both faithful and location-aware, effectively arranging scenes in a coherent manner when prompts contain multiple entities. Further details and visualization can be found in Supp. C.6.